OpenStudy (anonymous):

4c^2-12c+9 factor completely

OpenStudy (anonymous):

(3-2x)^2 or (2x-3)^2

OpenStudy (anonymous):

but how do you do it

OpenStudy (anonymous):

it's simple formula (a-b)^2 = a^2-2ab+b^2

OpenStudy (anonymous):

4c^2 -12c+9 multiply 9 x 4 =36 so two numbers that when u multiply equals 36 and add = -12. so the numbers are -6 and -6. so then u have 4c^2-6c-6c+9 then put them apart like (4c^2-6c) ( -6c +9) then factor out the two equations like 2c( 2c -3) -3 (2c -3) . finally take (2c-3) (2c-3)

OpenStudy (anonymous):

get it

OpenStudy (anonymous):

not really lol